# Generalized Heisenberg Dynamics Revisited

### v1

@@ 1 Introduction
Today, quantum mechanics and its extension to field theory, known as quantum field theory (collectively referred to here as quantum theory), are widely utilized across various areas of physics and have achieved remarkable success. In light of such success, there naturally arises an interest in exploring the foundations of quantum theory, investigating the scope of its applicability, and examining the potential for its extension. Indeed, motivated by such considerations, attempts have been made to extend matrix mechanics, and its fundamental structure has been studied from an algebraic perspective [11, 12, 13, 14]. This generalized mechanics is regarded as an extension of matrix mechanics, as it constructs physical quantities using an extended version of matrices, starting from generalized relations inspired by Bohr’s frequency condition and Ritz’s combination rule (see Figure 1).
@@
Furthermore, this generalized matrix mechanics exhibits similarities to Nambu mechanics (Nambu dynamics) and can be regarded as a quantum version of Nambu mechanics. Here, Nambu mechanics is a generalization of Hamiltonian dynamics through the extension of phase space based on the Liouville theorem [15] (see Figure 1). Specifically, in Ref. [13], the relationship between Nambu mechanics based on canonical triplets and the dynamics involving extended matrices with three indices is discussed from the viewpoint of the correspondence principle. However, this analysis is somewhat ad hoc and remains insufficient. Moreover, it is unclear whether a similar correspondence holds in the general case of dynamics involving extended matrices with $n$ indices. In Ref. [14], the theoretical formulation of dynamics based on extended matrices with $n$ indices was explored, but explicit expressions of frequency-related equations in terms of the eigenvalues of the Hamiltonians were not provided. In other words, further clarification of the theoretical structure of dynamics based on extended matrices with $n$ indices remains an open problem.
In this study, we address the question: “What kind of theoretical structure emerges when the correspondence principle is applied to Nambu mechanics?”. Answering this question is of considerable physical significance. This is because such an investigation can shed light on the ‘quantum’ or discrete version of the Nambu bracket, namely the quantum analogue of the Jacobian, potentially leading to a formulation of quantum theory for extended objects222 See Refs. [16, 17] for a comprehensive review of some aspects of Nambu mechanics and its applications in string/M-theory.. Furthermore, obtaining more concrete expressions for frequencies is expected to provide a clearer understanding of the underlying theoretical structure of an extended mechanics.
@@
The outline of this paper is as follows. In the next section, we review the structure of Hamiltonian dynamics and demonstrate how matrix mechanics can be derived from it using the correspondence principle. In section 3, we construct a dynamical framework involving extended matrices with three indices, starting from Nambu mechanics based on canonical triplets and applying the correspondence principle. In section 4, we develop a dynamical theory involving extended matrices with $n$ indices and study the properties of $n$-fold commutators defined on generalized matrices. In the final section, we present our conclusions and discussions. A detailed derivation of the equations presented in section 4 is provided in the Appendix A.
@@ 2.1 Hamiltonian dynamics
where $\varepsilon$ is an infinitesimal constant and $G=G(q,p)$ is a generating function. Then, using the Jacobi identity:
@@
where $A^{\prime}=A+\delta_{G}A$. From eqs. (2) and (4), we find that the time evolution of variables is realized by the canonical transformation whose generation function is the Hamiltonian.
@@
[…]
@@
[…]
@@
[…]
@@ 2.2 Heisenberg dynamics
where ${\hbar(\equiv{h}/({2\pi}))}$ is the reduced Planck constant and $[A(t),H]_{m~{}\!\!n}$ is a commutator defined by
@@
where $\varepsilon$ is an infinitesimal constant and $G_{m~{}\!\!n}$ is a generator. Then, using the Jacobi identity:
@@ 2.2.1 Fermionic harmonic oscillator
$H_{m~{}\!\!n}=\hbar\omega[C^{\dagger}(t),C(t)]_{m~{}\!\!n}=\hbar\omega\left(\sum_{k=1}^{2}C^{\dagger}_{m~{}\!\!k}(t)C_{k~{}\!\!n}(t)-\frac{1}{2}\delta_{m~{}\!\!n}\right).$ (43)
@@ 3.1 Nambu dynamics
where $\varepsilon$ is an infinitesimal constant and $G_{1}=G_{1}(x,y,z)$ and $G_{2}=G_{2}(x,y,z)$ are the generating functions. Then, using the fundamental identity [19]:
@@
Let us carry out the generalized canonical transformation $(x,y,z)\to(\theta,J_{1},J_{2})$, in which $\theta$ becomes a cyclic coordinate, using a function $W_{(1)}=W_{(1)}(y,z,J_{1},J_{2})=W_{1}dW_{2}$. The canonical triplets are related to each other in the following way,
$xdy\wedge dz-H_{1}dH_{2}\wedge dt=\theta dJ_{1}\wedge dJ_{2}-H^{\prime}_{1}dH^{\prime}_{2}\wedge dt+dW_{1}\wedge dW_{2},$ (51)
where $H^{\prime}_{1}=H^{\prime}_{1}(J_{1},J_{2})$ and $H^{\prime}_{2}=H^{\prime}_{2}(J_{1},J_{2})$ are new Hamiltonians. We notice that $H^{\prime}_{1}$ and $H^{\prime}_{2}$ do not depend on $\theta$. New canonical variables $\theta$, $J_{1}$ and $J_{2}$ satisfy the Nambu equation:
@@
From eq. (51), $x$, $\theta$, $H_{1}$, $H_{2}$, $H^{\prime}_{1}$ and $H^{\prime}_{2}$ are given by
$x=\frac{\partial(W_{1},W_{2})}{\partial(y,z)},~{}~{}\theta=-\frac{\partial(W_{1},W_{2})}{\partial(J_{1},J_{2})},~{}~{}H_{1}dH_{2}=H^{\prime}_{1}dH^{\prime}_{2},$ (55)
respectively. The period $T$ is given by the relation:
[…]
where we use the relation $({\partial(y,z)}/{\partial(E_{1},E_{2})})({\partial(H_{1},H_{2})}/{\partial(y,z)})=1$ and $dH_{1}\wedge dH_{2}=dH^{\prime}_{1}\wedge dH^{\prime}_{2}=dE_{1}\wedge dE_{2}$. When $\nu$ is regarded as a frequency, the following relation is obtained
[…]
@@
[…]
@@
We study the relationship between Nambu dynamics and Hamiltonian dynamics. Since $H_{1}$ and $H_{2}$ conserve, a particle moves on the path as the intersection of two surfaces described by $H_{1}=E_{1}$ and $H_{2}=E_{2}$. If we carry on the generalized canonical transformation $(x,y,z)\to(x^{\prime},y^{\prime},z^{\prime})$ with $z^{\prime}=H_{2}$, the Nambu equation becomes as
$\frac{dx^{\prime}}{dt}=\frac{\partial H_{1}}{\partial y^{\prime}},~{}~{}\frac{dy^{\prime}}{dt}=-\frac{\partial H_{1}}{\partial x^{\prime}},~{}~{}\frac{dz^{\prime}}{dt}=0,$ (60)
and it turns out to be Hamilton’s equation of motion by identifying $x^{\prime}$, $y^{\prime}$ and $H_{1}$ as $q$, $p$ and $H$, respectively.
@@ 3.2 Generalized Heisenberg dynamics
Let us formulate an extension of Heisenberg dynamics from Nambu dynamics, using the correspondence principle (see Figure 2).
@@
We assume that there is a one-to-one correspondence between values of $E_{1}(J_{1},J_{2})$ and $E_{2}(J_{1},J_{2})$ in Nambu dynamics and some discrete values $(E_{1})_{l~{}\!\!m}$ and $(E_{2})_{l~{}\!\!m}$ in its extension, and these dicrete values are anti-symmetric such as
@@
where we use the generalized rule (69). We refer to three-index variables, which follow the product rule (71), as cubic matrices333 Three-index objects have been introduced to construct a quantum version of the Nambu bracket in Ref. [21]. Their definition of the 3-fold product is different from ours..
@@
$=\frac{1}{i\hbar}\left[A(t),H_{1},H_{2}\right]_{l~{}\!\!m~{}\!\!n}(t),$ (75)
@@
as a counterpart of the Nambu equation (46). It is reconfirmed that the components with $l=m$, $m=n$ and/or $n=l$ on $A_{l~{}\!\!m~{}\!\!n}(t)$ are not independent of time, using eq. (77) and the relation $[B,C_{1},C_{2}]_{l~{}\!\!m~{}\!\!n}=0$ that holds concerning a cubic matrix $B$ with $B_{l~{}\!\!m~{}\!\!n}=0$ for $l\neq m$, $m\neq n$ and $n\neq l$ and arbitrary normal cubic matrices $C_{1}$ and $C_{2}$.
@@
with an infinitesimal constant $\varepsilon$ and normal cubic matrices $(G_{1})_{l~{}\!\!m~{}\!\!n}$ and $(G_{2})_{l~{}\!\!m~{}\!\!n}$, using a cubic matrix version of the fundamental identity555 When normal forms play the role of generators, the validity of the fundamental identity is confirmed for generalized matrices with $n(\geq 3)$ indices in subsection 4.2. In eq. (79), the Hamiltonians $H_{1}$ and $H_{2}$ act as generators.:
@@
it is shown that the equation is invariant such that
@@
$(G_{2})_{l~{}\!\!m~{}\!\!n}=(g_{2})_{l~{}\!\!m}(1-\delta_{l~{}\!\!n})\delta_{m~{}\!\!n}+(g_{2})_{m~{}\!\!n}(1-\delta_{m~{}\!\!l})\delta_{n~{}\!\!l}+(g_{2})_{n~{}\!\!l}(1-\delta_{n~{}\!\!m})\delta_{l~{}\!\!m}.$ (83)
From eqs. (77) and (78), we find that the time evolution of variables is generated by the transformation whose generators are the Hamiltonians.
Let us deal with eq. (77) as the master equation. Then, we find that there is another time-dependent solution such that

### v2

@@ 1 Introduction
Today, quantum mechanics and its extension to field theory, known as quantum field theory (collectively referred to here as quantum theory), are widely utilized across various areas of physics and have achieved remarkable success. In light of such success, there naturally arises an interest in exploring the foundations of quantum theory, investigating the scope of its applicability, and examining the potential for its extension. Indeed, motivated by such considerations, attempts have been made to extend matrix mechanics, and its fundamental structure has been studied from an algebraic perspective [11, 12, 13, 14]. This generalized mechanics is regarded as an extension of matrix mechanics, as it constructs physical quantities using an extended version of matrices, starting from generalized relations inspired by Bohr’s frequency condition and Ritz’s combination rule (see Figure 1). In Refs. [11, 12], taking Heisenberg’s matrix mechanics as a model, we constructed a dynamics based on generalized matrices with three indices, which we named cubic matrix mechanics. The equation of motion in this mechanics is formulated using a triple commutator, and since it involves only one counterpart of the Hamiltonian, it shares a structural similarity with Hamiltonian mechanics.
@@
Furthermore, in Refs. [13, 14], the framework was extended to include several counterparts of the Hamiltonian, and hence this generalized matrix mechanics exhibits similarities to Nambu mechanics (Nambu dynamics) and can be regarded as a quantum version of Nambu mechanics. Here, Nambu mechanics is a generalization of Hamiltonian dynamics through the extension of phase space based on the Liouville theorem [15] (see Figure 1). Specifically, in Ref. [13], the relationship between Nambu mechanics based on canonical triplets and the dynamics involving extended matrices with three indices is discussed from the viewpoint of the correspondence principle. However, this analysis is somewhat ad hoc and remains insufficient. Moreover, it is unclear whether a similar correspondence holds in the general case of dynamics involving extended matrices with $n$ indices. In Ref. [14], the theoretical formulation of dynamics based on extended matrices with $n$ indices was explored, but explicit expressions of frequency-related equations in terms of the eigenvalues of the Hamiltonians were not provided in a definite form such that they possess those classical counterparts. In other words, further clarification of the theoretical structure of dynamics based on extended matrices with $n$ indices remains an open problem.
In this study, we address the question: “What kind of theoretical structure emerges when the correspondence principle is applied to Nambu mechanics?”. Answering this question is of considerable physical significance. This is because such an investigation can shed light on the ‘quantum’ or discrete version of the Nambu bracket, namely the quantum analogue of the Jacobian, potentially leading to a formulation of quantum theory for extended objects [^2] See Refs. [16, 17] for a comprehensive review of some aspects of Nambu mechanics and its applications in string/M-theory.. Furthermore, obtaining more concrete expressions for frequencies is expected to provide a clearer understanding of the underlying theoretical structure of an extended mechanics.
@@
The outline of this paper is as follows. In the next section, we review the structure of Hamiltonian dynamics and demonstrate how matrix mechanics can be derived from it using the correspondence principle. In section 3, we construct a dynamical framework involving extended matrices with three indices, starting from Nambu mechanics based on canonical triplets and applying the correspondence principle. In section 4, we develop a dynamical theory involving extended matrices with $n$ indices and study the properties of $n$-fold commutators defined on generalized matrices. In the final section, we present our conclusions and discussions. A detailed derivation of the equations presented in section 4 is provided in the Appendix A. It is shown that the generalized Heisenberg equation can be reduced to Heisenberg’s equation of motion by taking a specific form of Hamiltonians in the Appendix B.
@@ 2.1 Hamiltonian dynamics
where $\varepsilon$ is an infinitesimal constant and $G=G(q,p)$ is a generating function. Then, using the derivation rule:
$\frac{d}{dt}\{A,B\}_{\rm PB}=\left\{\frac{dA}{dt},B\right\}_{\rm PB}+\left\{A,\frac{dB}{dt}\right\}_{\rm PB}$ (5)
and the Jacobi identity:
@@
where $A^{\prime}=A+\delta_{G}A$. For completeness, the derivation of eq. (7) is given as follows,
$\frac{dA^{\prime}}{dt}$ $=$ $\frac{d}{dt}\left(A+\delta_{G}A\right)=\frac{d}{dt}\left(A+\varepsilon\{A,G\}_{\rm PB}\right)$ (8)
$=$ $\frac{dA}{dt}+\varepsilon\left(\left\{\frac{dA}{dt},G\right\}_{\rm PB}+\left\{A,\frac{dG}{dt}\right\}_{\rm PB}\right)$
$=$ $\{A,H\}_{\rm PB}+\varepsilon\left(\{\{A,H\}_{\rm PB},G\}_{\rm PB}+\{A,\{G,H\}_{\rm PB}\}_{\rm PB}\right)$
$=$ $\{A,H\}_{\rm PB}+\varepsilon\{\{A,G\}_{\rm PB},H\}_{\rm PB}=\{A+\varepsilon\{A,G\}_{\rm PB},H\}_{\rm PB}$
$=$ $\{A+\delta_{G}A,H\}_{\rm PB}=\{A^{\prime},H\}_{\rm PB}.$
From eqs. (2) and (4), we find that the time evolution of variables is realized by the canonical transformation whose generation function is the Hamiltonian.
@@
Let us clarify the physical meanings of $\theta$, $\nu$ and $J$. We require that $\theta$ should satisfy ${\oint d\theta=1}$ during one cycle, and then $\nu$ is regarded as a frequency. Furthermore, we find that $J$ is identified with the area on the phase space:
$J=\oint_{\rm C}pdq,$ (14)
@@
$\oint d\theta=\oint d\left(\frac{\partial W}{\partial J}\right)=\frac{d}{dJ}\oint dW=\frac{d}{dJ}\oint_{\rm C}\frac{\partial W}{\partial q}dq=\frac{d}{dJ}\oint_{\rm C}pdq=1,$ (15)
@@
$\frac{dJ}{dE}=\oint_{\rm C}\frac{\partial p(q,E)}{\partial E}dq=\oint_{\rm C}\frac{1}{{\partial H}/{\partial p}}dq=\oint_{\rm C}\frac{1}{{dq}/{dt}}dq=\int_{0}^{T}dt=T,$ (16)
@@ 2.2 Heisenberg dynamics
where ${\hbar(\equiv{h}/({2\pi}))}$ is the reduced Planck constant and $A(t),H]_{m~\!\!n}$ is a commutator defined by
@@
where $\varepsilon$ is an infinitesimal constant and $G_{m~\!\!n}$ is a generator. Then, in a similar way to eqs. (5)–(8), using the derivation rule:
$\frac{d}{dt}[A,B]_{m~\!\!n}=\left[\frac{dA}{dt},B\right]_{m~\!\!n}+\left[A,\frac{dB}{dt}\right]_{m~\!\!n}$ (33)
and the Jacobi identity:
@@ 2.2.1 Fermionic harmonic oscillator
$H_{m~\!\!n}=\frac{1}{2}\hbar\omega[C^{\dagger}(t),C(t)]_{m~\!\!n}=\hbar\omega\left(\sum_{k=1}^{2}C^{\dagger}_{m~\!\!k}(t)C_{k~\!\!n}(t)-\frac{1}{2}\delta_{m~\!\!n}\right).$ (46)
@@ 3.1 Nambu dynamics
where $\varepsilon$ is an infinitesimal constant and $G_{1}=G_{1}(x,y,z)$ and $G_{2}=G_{2}(x,y,z)$ are the generating functions. Then, in a similar way to eqs. (5)–(8), using the derivation rule:
$\frac{d}{dt}\{A,B,C\}_{\rm NB}=\left\{\frac{dA}{dt},B,C\right\}_{\rm NB}+\left\{A,\frac{dB}{dt},C\right\}_{\rm NB}+\left\{A,B,\frac{dC}{dt}\right\}_{\rm NB}$ (52)
and the fundamental identity [19]:
@@
Here, we study the relationship between Nambu dynamics and Hamiltonian dynamics. Since $H_{1}$ and $H_{2}$ conserve, a particle moves on the path as the intersection of two surfaces described by $H_{1}=E_{1}$ and $H_{2}=E_{2}$. If we choose $z=H_{2}$, the Nambu equation becomes as
$\frac{dx}{dt}=\frac{\partial H_{1}}{\partial y},~~\frac{dy}{dt}=-\frac{\partial H_{1}}{\partial x},~~\frac{dz}{dt}=0,$ (55)
and it turns out to be Hamilton’s canonical equations of motion by identifying $x$, $y$ and $H_{1}$ as $q$, $p$ and $H$, respectively.
Let us consider a specific system with variables $\theta$, $J_{1}$ and $J_{2}$, in which they follow the Nambu equation and $\theta$ becomes a cyclic coordinate. In this system, as Hamiltonians $H^{\prime}_{1}$ and $H^{\prime}_{2}$ do not depend on $\theta$, the triplet satisfies the equation:
@@
If two systems described by eqs. (48) and (56) are related to the generalized canonical transformation $(x,y,z)\to(\theta,J_{1},J_{2})$, $\partial(\theta,J_{1},J_{2})/\partial(x,y,z)=1$ holds and the canonical triplets are related to each other in the following way,
$xdy\wedge dz-H_{1}dH_{2}\wedge dt=\theta dJ_{1}\wedge dJ_{2}-H^{\prime}_{1}dH^{\prime}_{2}\wedge dt+dW_{(1)},$ (59)
where $W_{(1)}$ is a 1-form. The period $T$ is given by the relation:
$T=\int_{0}^{T}dt=\oint_{\rm C}\frac{1}{{dx}/{dt}}dx=\oint_{\rm C}\frac{1}{{\partial(H_{1},H_{2})}/{\partial(y,z)}}dx=\oint_{\rm C}\frac{\partial(y,z)}{\partial(E_{1},E_{2})}dx,$ (60)
where C is a closed loop on the three-dimensional phase space, and we use the relation $({\partial(y,z)}/{\partial(E_{1},E_{2})})({\partial(H_{1},H_{2})}/{\partial(y,z)})=1$ and $dH_{1}\wedge dH_{2}=dH^{\prime}_{1}\wedge dH^{\prime}_{2}=dE_{1}\wedge dE_{2}$. When $\nu$ is regarded as a frequency, the following relation is obtained
$T=\frac{1}{\nu}=\frac{\partial(J_{1},J_{2})}{\partial(E_{1},E_{2})}=\oint_{\rm C}\frac{\partial(y,z)}{\partial(E_{1},E_{2})}dx,$ (61)
@@
$\iint\frac{dE_{1}\wedge dE_{2}}{\nu}=\iint dJ_{1}\wedge dJ_{2}=\iiint dy\wedge dz\wedge dx.$ (62)
@@
We have employed a generalized version of the Hamilton-Jacobi method for a system involving a cyclic coordinate; however, the applicability of our method is limited to specific classes of systems. Here, as an example, we consider the system described by the Hamiltonians:
$H_{1}=(a_{1}x+b_{1}y)^{2}+c_{1}z^{2}+f(a_{2}x+b_{2}y+z),~~H_{2}=a_{2}x+b_{2}y+z,$ (64)
where $a_{1}$, $b_{1}$, $c_{1}$, $a_{2}$ and $b_{2}$ are constants with $a_{1}b_{2}-a_{2}b_{1}>0$ and $c_{1}>0$. After the generalized canonical transformation such as
$x^{\prime}=\frac{1}{\sqrt{a_{1}b_{2}-a_{2}b_{1}}}(a_{1}x+b_{1}y),~~y^{\prime}=\frac{1}{\sqrt{a_{1}b_{2}-a_{2}b_{1}}}(a_{2}x+b_{2}y),~~z^{\prime}=a_{2}x+b_{2}y+z,$ (65)
the above Hamiltonians are written by
$H_{1}=(a_{1}b_{2}-a_{2}b_{1})x^{\prime 2}+c_{1}(a_{1}b_{2}-a_{2}b_{1})\left(\frac{z^{\prime}}{\sqrt{a_{1}b_{2}-a_{2}b_{1}}}-y^{\prime}\right)^{2}+f(z^{\prime}),~~H_{2}=z^{\prime},$ (66)
and this system is equivalent to one-dimensional harmonic oscillator system. In fact, $J_{1}$ and $J_{2}$ are expressed as
$J_{1}=\frac{\pi(E_{1}-f(J_{2}))}{\sqrt{c_{1}}(a_{1}b_{2}-a_{2}b_{1})},~~J_{2}=E_{2}$ (67)
and $\nu$ is given by
$\nu=\frac{\partial(E_{1},E_{2})}{\partial(J_{1},J_{2})}=\frac{\partial E_{1}}{\partial J_{1}}=\frac{\sqrt{c_{1}}(a_{1}b_{2}-a_{2}b_{1})}{\pi}$ (68)
where $E_{1}$ and $E_{2}$ are the values of $H_{1}$ and $H_{2}$, respectively. Incidentally, a generalization of the Hamilton-Jacobi theory that can be applied to any Nambu mechanical system has been formulated elegantly in Ref. [17] using the mathematical concept of fiber bundles.
@@ 3.2 Generalized Heisenberg dynamics
Let us formulate an extension of Heisenberg dynamics from a specific class of Nambu dynamics described by eq. (56), using the correspondence principle (see Figure 2).
@@
With the system described by eq. (56) in mind, we assume that there is a one-to-one correspondence between values of $E_{1}(J_{1},J_{2})$ and $E_{2}(J_{1},J_{2})$ in Nambu dynamics and some discrete values $(E_{1})_{l~\!\!m}$ and $(E_{2})_{l~\!\!m}$ in its extension, and these discrete values are anti-symmetric such as
@@
where we use the generalized rule (77). We refer to three-index variables, which follow the product rule (79), as cubic matrices [^3] Three-index objects have been introduced to construct a quantum version of the Nambu bracket in Ref. [21]. Their definition of the 3-fold product is different from ours..
@@
$=\frac{1}{i\hbar}\left[A(t),H_{1},H_{2}\right]_{l~\!\!m~\!\!n},$ (83)
@@
In the derivation of eq. (83), we use the relations $(h_{1})_{l~\!m}(1-\delta_{l~\!m})=(h_{1})_{l~\!m}$ and $(h_{2})_{l~\!m}(1-\delta_{l~\!m})=(h_{2})_{l~\!m}$ derived from the anti-symmetric property $(h_{1})_{m~\!l}=-(h_{1})_{l~\!m}$ and $(h_{2})_{m~\!l}=-(h_{2})_{l~\!m}$ (see eq. (69)).
@@
as a counterpart of the Nambu equation (49). It is reconfirmed that the components with $l=m$, $m=n$ and/or $n=l$ on $A_{l~\!\!m~\!\!n}(t)$ are independent of time, using eq. (85) and the relation $B,C_{1},C_{2}]_{l~\!\!m~\!\!n}=0$ that holds concerning a cubic matrix $B$ with $B_{l~\!\!m~\!\!n}=0$ for $l\neq m$, $m\neq n$ and $n\neq l$ and arbitrary normal cubic matrices $C_{1}$ and $C_{2}$. The equation (85) coincides with the fundamental equation of the generalized cubic matrix mechanics obtained in Ref. [13] through its structural analogy with Nambu mechanics.
It is natural to impose the triplet of cubic matrices $X_{l~\!\!m~\!\!n}$, $Y_{l~\!\!m~\!\!n}$ and $Z_{l~\!\!m~\!\!n}$ on the commutation relations: $\left[X,Y,Z\right]_{l~\!\!m~\!\!n}=i\hbar I_{l~\!\!m~\!\!n}$ and others=0=0 (or anti-commutation relations for cubic matrices describing fermionic variables) where $I$ is a specific constant cubic matrix (given by eq. (101)). In our generalized Heisenberg dynamics, the objective is to determine the values of Hamiltonians, i.e., $(h_{1})_{l~\!\!m}$, $(h_{1})_{m~\!\!n}$, $(h_{1})_{n~\!\!l}$, $(h_{2})_{l~\!\!m}$, $(h_{2})_{m~\!\!n}$ and $(h_{2})_{n~\!\!l}$ in eqs. (81) and (82),taking the equation of motion (85) as the master equation. However, since there is currently no method analogous to a unitary transformation in matrix mechanics, these values must be obtained directly.
Concerning the relationship between symmetries and conservation laws, cubic matrices $C^{a}$ that commute with the Hamiltonians, i.e., $C^{a},H_{1},H_{2}]=0$, are identified as conserved quantities, as understood from the generalized Heisenberg equation (85). Furthermore, arbitrary constant normal matrices become conserved quantities, because the relation $N_{1},N_{2},N_{3}]=0$ holds for any normal cubic matrices $N_{1}$, $N_{2}$ and $N_{3}$.
@@
with an infinitesimal constant $\varepsilon$ and normal cubic matrices $(G_{1})_{l~\!\!m~\!\!n}$ and $(G_{2})_{l~\!\!m~\!\!n}$, using the derivation rule:
$\frac{d}{dt}[A,B,C]_{l~\!\!m~\!\!n}=\left[\frac{dA}{dt},B,C\right]_{l~\!\!m~\!\!n}+\left[A,\frac{dB}{dt},C\right]_{l~\!\!m~\!\!n}+\left[A,B,\frac{dC}{dt}\right]_{l~\!\!m~\!\!n}$ (87)
and a cubic matrix version of the fundamental identity [^5] When normal forms play the role of generators, the validity of the fundamental identity is confirmed for generalized matrices with $n(\geq 3)$ indices in subsection 4.2. In eq. (88), the Hamiltonians $H_{1}$ and $H_{2}$ act as generators.:
@@
it is shown that the equation (85) is invariant such that
@@
$(G_{2})_{l~\!\!m~\!\!n}=(g_{2})_{l~\!\!m}(1-\delta_{l~\!\!n})\delta_{m~\!\!n}+(g_{2})_{m~\!\!n}(1-\delta_{m~\!\!l})\delta_{n~\!\!l}+(g_{2})_{n~\!\!l}(1-\delta_{n~\!\!m})\delta_{l~\!\!m},$ (92)
where $(g_{1})_{m~\!l}$s and $(g_{2})_{m~\!l}$s are anti-symmetric constants. Eq. (89) is obtained in a similar way to the counterpart in the Nambu mechanics. From eqs. (85) and (86), we find that the time evolution of variables is generated by the transformation whose generators are the Hamiltonians.
When we deal with eq. (85) as the master equation, we find that various systems can be described as solutions of this equation. Here, we present another time-dependent solution satisfying the generalization of Ritz’s combination rule (77). It is written by
@@
The present solution agrees with the solution of the fundamental equation of cubic matrix mechanics derived in Refs. [11, 12]. In the Appendix B, it is shown that the generalized matrix mechanics can describe physical systems in conventional matrix mechanics.